Convert the following wavelengths to frequencies (Hz):

a. 0.71 μm

b. 31.20 μm

c. 419 nm

d. 1385 nm

e. 17.31 cm 


1
Expert's answer
2021-10-14T11:31:15-0400

"\lambda = \frac{v}{f}"

"f = \frac{v}{\lambda}"

"\text{Since no propagation medium is specified, the }"

"\text{default is an electromagnetic wave in a vacuum.}"

"c \approx 3*10^8\frac{m}{c}"

"f= \frac{c}{\lambda}"


"a. \lambda= 0.71 μm = 0.71*10^{-6}m"

"f= \frac{c}{\lambda}=\frac{3*10^8}{0.71*10^{-6}}=4.23*10^{14}Hz"


"b. \lambda = 31.20 μm = 31.2*10^{-6}m"

"f= \frac{c}{\lambda}=\frac{3*10^8}{31.2*10^{-6}}=9.62*10^{12}Hz"


"c. \lambda= 419 nm= 419*10^{-9}m"

"f= \frac{c}{\lambda}=\frac{3*10^8}{419*10^{-9}}=7.16*10^{14}Hz"


"d. \lambda= 1385 nm= 1385*10^{-9}m"

"f= \frac{c}{\lambda}=\frac{3*10^8}{1385*10^{-9}}=2.17*10^{14}Hz"


"e. \lambda=17.31 cm =17.31*10^{-2}m"

"f= \frac{c}{\lambda}=\frac{3*10^8}{17.31*10^{-2}}=1.73*10^{9}Hz"